3. (20) In how many ways can the word KUTKUT be arranged so that no two identical letters are next to each other?
Answer #1

= 5! Now, KUTKUT is a 6 letter world with qks, a Ts and 2Us. KIT, U are So, Total no. of ways. repeating two 2!2!2! timesnoi of ways noi of ways when aks 20s are together and TIT are and ther 2 world [KK QU TT So, no. of ways = 4! lly no. of wa= 6! 2121 21 – 3,5! 21 21 + 3.41 21 – 3) 6x5X 4x3x2 2x2x2 – 3xSxyx3x2 + 3x4x342-6. xxx So o = 96 - 96 + 36-6 = 30 no. of w no